# A Survey of Language-Based Communication in Robotics

### v1

@@ Abstract.
Embodied robots which can interact with their environment and neighbours are increasingly being used as a test case to develop Artificial Intelligence. This creates a need for multimodal robot controllers which can operate across different types of information including text. Large Language Models are able to process and generate textual as well as audiovisual data and, more recently, robot actions. Language Models are increasingly being applied to robotic systems; these Language-Based robots leverage the power of language models in a variety of ways. Additionally, the use of language opens up multiple forms of information exchange between members of a human-robot team. This survey motivates the use of language models in robotics, and then delineates works based on the part of the overall control flow in which language is incorporated. Language can be used by human to task a robot, by a robot to inform a human, between robots as a human-like communication medium, and internally for a robot’s planning and control. Applications of language-based robots are explored, and finally numerous limitations and challenges are discussed to provide a summary of the development needed for language-based robotics moving forward. Links to each paper and, if available, source code are made available in the accompanying site [^1] Website with table providing links to papers and code.
1. Introduction
@@
A popular trend in Artificial Intelligence is toward powerful, multimodal models that operate on multiple domains of input and output (Reed et al., 2022). This often centres around foundational models based on vision and language (Di Palo et al., 2023). This is especially true in the field of robotics, where future robotic systems could operate on a single architecture which combines learning, understanding, and actions across different modalities (Driess et al., 2023). Core to this approach is the integration of language models; not only does this potentially improve the power and potential of robotic systems, but it also introduces many promising interaction opportunities. Language can be included in many different possible elements of the total architecture, each providing different advantages.
@@
Figure 1. Overview of this survey. Works are categorised by direction of communication, and advantages, applications, and limitations are discussed.
@@
Large Language Models (LLMs) are a rapidly advancing technology that show promise in many diverse areas of science and engineering (Yang et al., 2023a). Among the many potential advantages of a language-based AI landscape is that the natural languages encapsulate deep meaning and complex relationships between semantic concepts; a sentence of only a few words may represent a detailed idea, and minor alterations can substantially alter the meaning (Crocker et al., 2016). Virtually all human interaction, coordination, and problem solving is in some way facilitated by language, and yet robots have traditionally never benefited from this. Despite existing to replace human workers, robots in the past have generally been controlled though direct teleoperation, predefined paths, or reinforcement learning (Kober et al., 2013). An emerging trend in robotics is the integration of LLMs into the robotic command and control structure (Zeng et al., 2023). This development goes beyond the simple paradigm of labelling predefined, hand-written behaviours which can be commanded by a human user, and instead treats language as an input space like any other. Language models instead utilise the concept of “embeddings” to capture the semantic meaning of a piece of text (Reimers and Gurevych, 2019). This allows them to operate more freely and flexibly with respect to whatever wording or structure of textual input is provided. This also presents an opportunity for greater generalisation, where the language model can learn from examples in manner that need not conform with a given structure.
Kim et al. explored the use of LLMs in robots by delineating communication, planning, perception, and control. Their study of communication covered language generation and understanding including inter-robot and human communication but in limited depth. Both Guo et al. and Wang et al. have surveyed LLMs in multi-agent systems, but neither covers human interaction in significant detail, and neither focuses on robotics. Zeng et al. covered perception, decision-making, control, and interaction and include an excellent glossary of terms and models; they do not, however, discuss the multi-agent aspect of LLMs in robots. Zhang et al. provided a detailed exploration of human-robot interaction via physical, teleoperation, and conversation including language, and Atuhurra covered multiple applications and architectures for LLMs as well as societal norms surrounding them. None of these works fully explores the entire context of a mixed team of humans and robots; in this scenario language can be employed in every part of the system. This survey aims to fill this gap.
@@
Language presents several opportunities in robotics, both from a research perspective and one of future systems deployed in the real world (section 2). We propose that any system with one or many robots and one or many humans working alongside each other can be considered in a manner where language can be inserted as a component of any interface between them (section 3). Language can now be used as an input space to robotic models so they can respond to hand-written instructions from humans without the need to conform to a rigid format or grammar (section 3.1). Robots can also generate language that is used to explain its decisions and behaviour to a human (section 3.2). This can be further extended to have robots communicate with each other using language as both an input and output, simulating human collectives who plan and coordinate in this way (section 3.3). Emerging research integrates language internally into the model as part of its architecture (section 3.3). We describe these in detail and then summarise the potential future applications of this concept (section 4), as well as the current limitations (section 5).
2. Why Use Language?
@@
The scope of research using LLMs has become extremely broad. There are being applied to a wide variety of domains outside of standard text generation including education (Kasneci et al., 2023), medicine (Thirunavukarasu et al., 2023; Liévin et al., 2024), programming (Xu et al., 2022), and video games (Raad et al., 2024). They are also beginning to be employed in multi-agent systems (Guo et al., 2024; Wang et al., 2024c), human-robot interaction (Zhang et al., 2023a; Atuhurra, 2024) and robotics generally (Kim et al., 2024a; Zeng et al., 2023). This section suggests four core reasons to investigate LLMs in robotics.
2.1. Knowledge and Inference
@@
Language models are trained on a large and varied corpus of written text, online web data, and community-maintained repositories of information (Minaee et al., 2024; Liu et al., 2024a). As such, this knowledge is distilled into the models, granting them an extremely broad and often quite deep ‘understanding’ (Hinton, 2023). While the true nature of their ability to reason is debated (Cohen, 2024; Browning and LeCun, 2023; Häggström, 2023), they seem capable of combining knowledge and ideas together to produce new outputs (Romera-Paredes et al., 2024). This makes them well suited to interpreting the true meaning behind the tasks they are set, as well as being able to expand on observations and details by combining learned knowledge in order to achieve their goals.
2.2. Zero-shot Adaptability
@@
Language models have been proposed as excellent few-shot or even zero-shot learners (Brown et al., 2020). In robotics the environment is unpredictable and unforeseen elements are common. By applying language models to the problem, we can benefit from their ability to guess about the impacts of novel events. For example, a robot may never have trained for a particular terrain or weather, but language can offer an estimation of its impact (Shek et al., 2023). Similarly, new tasks that have not been trained for may be decomposed by a language model into performable substeps, or else the model may prove effective at understanding which parts of a complex task are important to consider and how to handle them. The ability for language models to perform in-context learning may be a useful tool in empowering other approaches, even if not to replace them entirely (Dasgupta et al., 2023).
2.3. Interactable and Understandable
@@
Language is the fundamental means of direct interaction between humans. Both structured and unstructured collectives generally utilise language to coordinate, plan, and exchange information (Mirsky et al., 2020). To understand and communicate often requires a dialogical component in order to ground information and establish a shared knowledge and perception of the world (Clark and Brennan, 1991). As an extremely dense embedding of information, a few words can encapsulate deep meaning. The specific framing and interplay between the words allows a listener to comprehend a complex, high-level idea from a simple sentence (Minsky, 1988). Allowing robots to use language opens up their command, control, and even reasoning to human intervention; there is no requirement for operators and bystanders to understand programming languages or technical commands. By building robots which can emulate human communication, they can be commanded by humans or can explain their actions in the same way a human worker might. They can also cooperate with each other in an understandable way (Stone et al., 2010). A language-based architecture, or one where language is used to interface between parts of a policy or agents in a team allows robots to be swapped-out for humans at any level, and also allows humans to understand and assist robots easily. This also makes it much easier for subject matter experts to work alongside robots without technical knowledge of their processes.
2.4. Biological Inspiration and Language-as-Intelligence
@@
Much of the development of Artificial Intelligence has been ‘bottom-up’ –– taking inspiration from biological intelligence and building computer systems which model the same structures and processes in the digital realm (Frenkel et al., 2021). Neural networks are a common example of this, where relatively naïve approaches allow us to emulate the way humans learn; as nature has arrived at a formula for human intelligence this surely must be one means of achieving it digitally (if not necessarily the best (Hinton, 2023)). Rodney Brooks postulated that embodiment was key to intelligence, making robots a good test case for building AI (Brooks, 1991), while others also argue that language is in some sense core to human thinking and reasoning (Goldstein and Papert, 1977; Bloom and Keil, 2001; Gleitman and Papafragou, 2013; Pinker, 2003), even perhaps that language is a “system of thought” sui generis (Chomsky, 2011), or at least shapes the way we perceive and experience the world around us (Boroditsky, 2011). It may be the case that empowering physical agents with the ability to interpret, internally reason with, and output language could substantively improve their intelligence. Regardless, language as a planning and communication tool naturally mimics human intelligence and behaviour in many settings, so it natural to explore in the context of robots as human-like workers.
3. A Future Paradigm for Multi-Robot Systems
@@
3.1. Human to Robot Communication
@@
The most suited use-case of language models in robotics is the direct commanding of robots. Classical approaches have defined a library of skills and commands with textual labels that can be used to control a robot (Abioye et al., 2018). However the introduction of LLMs redefines this paradigm; what was once a discrete input space defined by a programmer can now be learned at a higher level. By considering a textual (or, by extension, visual or gestural) command as encoding some meaning which exists in an embedding space, it can be theorised that a model can interpret an instruction never before seen, and internally relate it to previous inputs. This is encapsulated by the fundamental concept of machine learning — generalisation, but it is an important and nontrivial step to move from generalising to classify unseen images to interesting and acting based on the semantic meaning of a sentence. The “Human to Robot Communication” category describes work which focuses on using human-style commands that may otherwise be issued to other humans, such as “pick up the red ball” and creating models that map this to an action.
@@
Table 1. A selection of works using language for Human-to-Robot tasking. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Italics indicate a simulator, and “language used to” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Robot/sim Language used to
Task Breakdown Dorbala et al. GPT-3 RobotTHOR Generate language-based plans
Brohan et al. GPT-3.5 Franka Robot Generate plans from commands
Wu et al. TaPA AI2THOR Generate plans from commands
Kakodkar et al. GPT-3 Clearpath Jackal UGV Generate plans from commands
Huang et al. GPT-3 VirtualHome Generate plans from commands
Rana et al. GPT-4 Robot arm on wheels Plan large tasks
Kapelyukh et al. DALL-E Robot arm Generate target state
Wang et al. GPT-4 Ravens Generate sim data
Zhi-Xuan et al. GPT-3 VirtualHome Determine human intentions
Code and Rewards Singh et al. GPT-3 Virtual Home & Franka Panda Write code from a library
Hu et al. Various RoboEval Write code from a library
Lin et al. GPT-3.5 Franka Panda Write code from a library
Huang et al. GPT-3 AI2THOR Write code from a library
Liang et al. GPT-3 UR5e robot arm Write code from a library
Lin et al. GPT-3.5 Franka Panda Write code from list of skills
Yu et al. GPT-4 MuJuCo Generate reward function
Hu and Sadigh GPT-3.5 N/A Generate reward function
Zhang et al. GPT-2 N/A Write code from descriptions
Ma et al. GPT-4 IsaacGym Writing and updating code
Narin GPT-4V IsaacGym Writing and updating code
Yu et al. GPT-3 Robot arm Generate training examples
Wu et al. PaLM 540B Kinova Gen3 7-DoF Write plans in code
Integrated Input Wen et al. GPT-3.5 Franka Robot Augmented into the model
Shah et al. GPT-3 Clearpath Jackal UGV Augmented into input space
Ha et al. GPT-3 & LLAMA2 MuJoCo Write success labelling function
Zhu et al. GPT-3.5 N/A Integrated into the model
Bärmann et al. GPT-3.5 ARMAR-6 Integrated as memory and to adapt
Liu et al. GPT-2 Overcooked Integrated into the model
Bing et al. GloVe Metaworld and robot arm Integrated into MDP
Ni et al. INSTRUCTOR PUDUbot2 Integrated into the model
Yow et al. S-BERT xArm-6 manipulator Integrated into the model
Bucker et al. BERT Panda robot arm Embedding input into architecture
Lynch et al. LCBC [^2] Language conditioned behavioral cloning Language-Table and xArm6 Integrated into the model
Cui et al. GPT-3 Franka Emika Panda Modify policy with embedding
Huang et al. GPT-4 Franka Emika Panda Modify policy with embedding
Karamcheti et al. Distil-RoBERTa Franka Panda Disambiguate manual input
Gu et al. RT-Trajectory Everyday Robots arm Generate trajectories
Shi et al. GPT-4V ALOHA Modify robot behaviour
3.1.1. Task Breakdown and Planning

### v3

@@ Abstract
Embodied robots which can interact with their environment and neighbours are increasingly being used as a test case to develop Artificial Intelligence. This creates a need for multimodal robot controllers that can operate across different types of information, including text. Large Language Models are able to process and generate textual as well as audiovisual data and, more recently, robot actions. Language Models are increasingly being applied to robotic systems; these Language-Based robots leverage the power of language models in a variety of ways. Additionally, the use of language opens up multiple forms of information exchange between members of a human-robot team. This survey motivates the use of language models in robotics, and then delineates works based on the part of the overall control flow in which language is incorporated. Language can be used by a human to task a robot, by a robot to inform a human, between robots as a human-like communication medium, and internally for a robot’s planning and control. Applications of language-based robots are explored, and numerous limitations and challenges are discussed to provide a summary of the development needed for the future of language-based robotics. [^1] Links to each paper and, if available, source code are provided at the accompanying webpage: https://sooratilab.com/publications/2024/A-Survey-of-Language-BasedCommunication-in-Robotics.php.
1 Introduction
@@
A popular trend in Artificial Intelligence is toward powerful, multimodal models that operate on multiple domains of input and output Reed et al., (2022). This often centres around foundational models based on vision and language Di Palo et al., (2023). This is especially true in the field of robotics, where future robotic systems could operate on a single architecture that combines learning, understanding, and actions in different modalities Driess et al., (2023). Core to this approach is the integration of language models; not only does this potentially improve the power and potential of robotic systems, but it also introduces many promising interaction opportunities. Language can be included in various elements of the total architecture, each providing different advantages.
@@
Figure 1: Overview of this survey. Works are categorised by direction of communication between members of the human-robot team. Advantages, applications, and limitations are also discussed.
@@
Large Language Models (LLMs) are a rapidly advancing technology that shows promise in many diverse areas of science and engineering Yang et al., 2023a . Among the many potential advantages of a language-based AI landscape is that natural language encapsulates deep meaning and complex relationships between semantic concepts; a sentence of only a few words may represent a detailed idea, and minor alterations can substantially alter the meaning Crocker et al., (2016). Virtually all human interaction, coordination, and problem-solving is in some way facilitated by language, and yet robots have traditionally never benefited from this. Despite being intended to replace human workers, robots in the past have generally been controlled through direct teleoperation, predefined paths, or reinforcement learning Kober et al., (2013). An emerging trend in robotics is the integration of LLMs into the robotic command and control structure Zeng et al., (2023). This development goes beyond the simple paradigm of labelling predefined, hand-written behaviours which can be commanded by a human user, and instead treats language as an input space like any other. Language models instead utilise the concept of “embeddings” to capture the semantic meaning of a piece of text Reimers and Gurevych, (2019). This allows them to operate more freely and flexibly with respect to whatever wording or structure of textual input is provided. It also presents an opportunity for greater generalisation, where the language model can learn from examples in a manner that does not need to conform to a given structure.
Kim et al., 2024b explored the use of LLMs in robots by dividing them based on the application of the LLM; grouping them into communication, planning, perception, and control methods. Their study of communication covered language generation and understanding including some human-interaction aspects. This survey offers a review of the various parts of a robot or robots on which language can be applied. Wang et al., 2024b focus only on the single-agent case, they explore LLMs for agent memory as well as other parts of the architecture of an agent into which language can be integrated. Guo et al., (2024) reviewed LLMs for multi-agent systems, exploring different applications of the technology and the various architecture designs. Each of these works covers the agent-based model of LLMs very well, although both are outside the scope of robotics. Zeng et al., (2023), like Kim et al., 2024b , divided works by application in robotics into the areas of perception, decision-making, control, and interaction. They also include a glossary of terms and models that summarises the key ideas and models for LLM-based robots. Zhang et al., 2023a provided an exploration of human-robot interaction by categorising the use of LLMs for physical, teleoperation, and conversation methods. Atuhurra, (2024) covered multiple applications and architectures for LLMs as well as societal norms surrounding them. Both of these works are highly relevant to the relationship between a language-based robot and its operators.
None of the previous surveys explore the entire context of a mixed team of humans and robots. A scenario where one or more humans must control or operate alongside a team of multiple agents demands an understanding of how language could be employed in every part of the system. Viewing the team of robots and operators as a collective who communicate, report, and control each other and themselves through the use of language provides an appropriate grouping of works based on where in the system they could feasibly be deployed. This survey aims to fill this gap by classifying and exploring works based on the part of the overall team structure in which language can be integrated
@@
Language presents several opportunities in robotics, both from a research perspective and one of the future systems deployed in the real world (section 2). We propose that any system with one or many robots and one or many humans working alongside each other can be considered in a manner where language can be inserted as a component of any interface between them (section 3). Language can now be used as an input space for robotic models so they can respond to hand-written instructions from humans without the need to conform to a rigid format or grammar (section 3.1). Robots can also generate language that is used to explain or validate their decisions and behaviour to a human (section 3.2). This can be further extended to have robots communicate with each other using language as both an input and output, simulating human collectives who plan and coordinate in this way (section 3.3). Emerging research integrates language internally into the model as part of its architecture (section 3.4). We describe these in detail and then summarise some potential future applications of this concept (section 4), as well some key limitations and challenges in this area (section 5).
2 Why Use Language?
@@
The scope of research using LLMs has become extremely broad. There are being applied to a wide variety of domains outside of standard text generation including education Kasneci et al., (2023), medicine Thirunavukarasu et al., (2023); Liévin et al., (2024), programming Xu et al., (2022), and video games Raad et al., (2024). They are also beginning to be employed in multi-agent systems Guo et al., (2024); Wang et al., 2024b , human-robot interaction Zhang et al., 2023a ; Atuhurra, (2024) and robotics generally Kim et al., 2024b ; Zeng et al., (2023). This section suggests four core reasons to investigate LLMs in robotics.
2.1 Knowledge and Inference
@@
Language models are trained on a large and varied corpus of written text, online web data, and community-maintained repositories of information Minaee et al., (2024); Liu et al., 2024b . As such, this knowledge is distilled into the models, granting them an extremely broad and often quite deep ‘understanding’ Hinton, (2023). While the true nature of their ability to reason is debated Cohen, (2024); Browning and LeCun, (2023); Häggström, (2023), they seem capable of combining knowledge and ideas together to produce new outputs Romera-Paredes et al., (2024). This makes them well suited to interpreting the true meaning behind the tasks they are set, as well as being able to expand on observations and details by combining learned knowledge in order to achieve their goals.
2.2 Zero-shot Adaptability
@@
Language models have been proposed as excellent few-shot or even zero-shot learners Brown et al., (2020). In robotics the environment is unpredictable and unforeseen elements are common. By applying language models to the problem, we can benefit from their ability to guess about the impacts of novel events. For example, a robot may never have trained for a particular terrain or weather, but language can offer an estimation of its impact Shek et al., (2023). Similarly, new tasks that have not been trained for may be decomposed by a language model into performable substeps, or else the model may prove effective at understanding which parts of a complex task are important to consider and how to handle them. The ability for language models to perform in-context learning may be a useful tool in empowering other approaches, even if not to replace them entirely Dasgupta et al., (2023).
2.3 Interactable and Understandable
@@
Language is the fundamental means of direct interaction between humans. Both structured and unstructured collectives generally utilise language to coordinate, plan, and exchange information Mirsky et al., (2020). To understand and communicate often requires a dialogical component in order to ground information and establish a shared knowledge and perception of the world Clark and Brennan, (1991). As an extremely dense embedding of information, a few words can encapsulate deep meaning. The specific framing and interplay between the words allows a listener to comprehend a complex, high-level idea from a simple sentence Minsky, (1988). Allowing robots to use language opens up their command, control, and even reasoning to human intervention; there is no requirement for operators and bystanders to understand programming languages or technical commands. By building robots which can emulate human communication, they can be commanded by humans or can explain their actions in the same way a human worker might. They can also cooperate with each other in an understandable way Stone et al., (2010). A language-based architecture, or one where language is used to interface between parts of a policy or agents in a team allows robots to be swapped-out for humans at any level and also allows humans to understand and assist robots easily. This also makes it much easier for subject matter experts to work alongside robots without technical knowledge of their processes.
2.4 Biological Inspiration and Language-as-Intelligence
@@
Much of the development of Artificial Intelligence has been ‘bottom-up’ –- taking inspiration from biological intelligence and building computer systems which model the same structures and processes in the digital realm Frenkel et al., (2021). Neural networks are a common example of this, where relatively naïve approaches allow us to emulate the way humans learn; as nature has arrived at a formula for human intelligence, this surely must be one means of achieving it digitally, even if not necessarily the best Hinton, (2023). Rodney Brooks postulated that embodiment was key to intelligence, making robots a good test case for building AI Brooks, (1991), while others also argue that language is in some sense core to human thinking and reasoning Goldstein and Papert, (1977); Bloom and Keil, (2001); Gleitman and Papafragou, (2013); Pinker, (2003), even perhaps that language is a “system of thought” sui generis Chomsky, (2011), or at least shapes the way we perceive and experience the world around us Boroditsky, (2011). It may be the case that empowering physical agents with the ability to interpret, internally reason with, and output language could substantively improve their intelligence. Regardless, language as a planning and communication tool naturally mimics human intelligence and behaviour in many settings, so it may be productive to explore in the context of robots as human-like workers.
3 The Paradigm for Communication in Multi-Robot Systems
@@
3.1 Human to Robot Communication
@@
The most suited use-case of language models in robotics is the direct commanding of robots. Classical approaches have defined a library of skills and commands with textual labels that can be used to control a robot Abioye et al., (2018). However, the introduction of LLMs redefines this paradigm; what was once a discrete input space defined by a programmer can now be learned at a higher level. By considering a textual (or, by extension, visual or gestural) command as encoding some meaning which exists in an embedding space, it can be theorised that a model can interpret an instruction never before seen and internally relate it to previous inputs. This is encapsulated by the fundamental concept of machine learning — generalisation, but it is an important and nontrivial step to move from generalising based on training to data, to understanding and acting based on the semantic meaning of a sentence. The “Human to Robot Communication” category describes work which focuses on using human-style commands that may otherwise be issued to other humans, such as “pick up the red ball” and creating models that map this to an action.
@@
Table 1: A selection of works using language for Human-to-Robot tasking. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). Italics indicate a simulator, and “Application” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Physical Robot / Simulation Application Task Breakdown Dorbala et al., (2023) GPT-3 RoboTHOR Generating plans Brohan et al., 2023b GPT-3.5 Franka Robot Generating plans Wu et al., 2023d TaPA AI2THOR Generating plans Kakodkar et al., (2023) GPT-3 Clearpath Jackal UGV Generating plans Huang et al., 2022a GPT-3 VirtualHome Generating plans Rana et al., (2023) GPT-4 Robot arm on wheels Generating plans Kapelyukh et al., (2023) DALL-E Robot arm Generating target state Wang et al., 2023d GPT-4 Ravens Generating simulation data Zhi-Xuan et al., (2024) GPT-3 VirtualHome Determining human intentions Code and Rewards Singh et al., (2023) GPT-3 Virtual Home & Franka Panda Writing code Hu et al., (2024) Various RoboEval Writing code Lin et al., 2023b GPT-3.5 Franka Panda Writing code Huang et al., 2023a GPT-3 AI2THOR Writing code Liang et al., (2023) GPT-3 UR5e robot arm Writing code Lin et al., 2023a GPT-3.5 Franka Panda Writing code Yu et al., 2023b GPT-4 MuJuCo Generating reward function Hu and Sadigh, (2023) GPT-3.5 N/A Generating reward function Zhang et al., 2023e GPT-2 N/A Writing code Ma et al., (2023) GPT-4 IsaacGym Writing code Narin, (2024) GPT-4V IsaacGym Writing code Yu et al., 2023a GPT-3 Robot arm Generating training examples Wu et al., 2023a PaLM 540B Kinova Gen3 7-DoF Writing code Integrated Input Wen et al., (2024) GPT-3.5 Franka Robot Augmenting into model Shah et al., (2023) GPT-3 Clearpath Jackal UGV Augmenting into input space Ha et al., (2023) GPT-3 & LLAMA2 MuJoCo Writing success labelling function Zhu et al., 2024a GPT-3.5 N/A Integrating into model Bärmann et al., (2023) GPT-3.5 ARMAR-6 Integrating as memory Liu et al., 2023a GPT-2 Overcooked Integrating into model Bing et al., (2023) GloVe Metaworld and robot arm Integrating into MDP Ni et al., (2023) INSTRUCTOR PUDUbot2 Integrating into model Yow et al., (2024) S-BERT xArm-6 manipulator Integrating into model Bucker et al., (2023) BERT Panda robot arm Embedding input into architecture Lynch et al., (2023) LCBC [^2] Language conditioned behavioral cloning Language-Table and xArm6 Integrating into model Cui et al., 2023b GPT-3 Franka Emika Panda Modifying with embedding Huang et al., 2023b GPT-4 Franka Emika Panda Modifying with embedding Karamcheti et al., (2022) Distil-RoBERTa Franka Panda Disambiguating manual input Gu et al., 2023a RT-Trajectory Everyday Robots arm Generating trajectories Shi et al., (2024) GPT-4V ALOHA Modifying behaviour
3.1.1 Task Breakdown and Planning